A pencil is knocked horizontally off the edge of a desk at 2.0 ms -1. The desk is 65 cm high.
a) How long does it take the pencil to reach the floor? Ignore air resistance

b) what horizontal distance does the pencil travel during this time?
1
Expert's answer
2019-10-24T09:34:53-0400

a)


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(0.65)}{9.8}}=0.36\ s"

b)


"x=vt=v\sqrt{\frac{2h}{g}}=2\sqrt{\frac{2(0.65)}{9.8}}=0.73\ m"